40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 489     Entry time: Tue May 20 18:33:01 2008
Author: Andrey, John 
Type: Configuration 
Category: IOO 
Subject: Mode Cleaner is locked again 

It was noticed by Mr.Adhikari earlier today that the MC became unlocked at about 11AM.

There is no clear understanding what caused the problem.

Trying to restore the modecleaner locking, we noticed with John that the beam was not centered on the wavesensors (WFS1. WFS2 on the screen "C1IOO_LockMC.adl"). We decided to adjust the beam position moving slightly the bias sliders for pitch and yaw degrees of freedom for MC1.
This allowed to make the MC locked.

Old positions for the MC1 sliders: Pitch = 2.9934, Yaw = -0.6168;
New positions --------//---------: Pitch = 3.0604, Yaw = -0.7258.

At the same time, FSS for PSL is still showing the values in the range 0.720 - 0.750 which is lower than the usual values. The indicator for FSS value is yellow when it is below 0.750.
ELOG V3.1.3-